A process occurs at a rate of 4.5 per time period?       What assumptions need to...

  1. A process occurs at a rate of 4.5 per time period?

  1.       What assumptions need to be made to model this process with a Poisson distribution?

         Applying the assumptions in part a), what is the probability that:

  1.       0 occurrences will appear in the next time period?
  2.       5 or fewer occurrences will appear in the next time period?
  3.       6 or more occurrences will appear in the next time period?
Answer #1

    Rate = 4.5 per time period                              
   Let X be the number of occurrence of the process                              
   Then X follows Poisson (λ = 4.5)                              
                                  
a)   Assumptions for modeling the process with Poisson distribution are                               
   1) The occurences of the process are independent.                               
   2) The average rate at which process occurs is constant.                              
   3) Two processes cannot occur at exactly the same instant                              
   4) The number of occurences of the process takes discrete values from 0, 1, 2, …                              
                                  
b)   To find P(0 occurences will appear in the next time period)                              
   that is to find P(X = 0)                              
e-222 P(X = 0) = e-4.5 (4.5) 0!

     P(X = 0) = 0.0111
   P(0 occurences will appear in the next time period) = 0.0111                              
                                  
                                  
c)   To find P(5 or fewer occurences will appear in the next time period)                              
   that is to find P(X ≤ 5)                              
                                  
   P(X ≤ 5) = P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3) + P(X = 4) + P(X = 5)                              
e-4.5(4.520e-4.5(4.5)1 e-4.5(4.5)2 e-4.5(4.5)3 = - +- - +- - + ~ み、 e-4.5 (4.504 e-4.5(4.5)5 5! ~ 1! 3!
                   = 0.7029                              
   P(5 or fewer occurences will appear in the next time period) = 0.7029                              
                                  
d)   To find P(6 or more occurences will appear in the next time period)                              
   that is to find P(X ≥ 6)                              
   P(X ≥ 6) = 1 - P(X < 6)                              
                    = 1 - P(X ≤ 5)                              
                    = 1 - 0.7029           …... As calculated in (c)                  
                    = 0.2971                              
   P(6 or more occurences will appear in the next time period) = 0.2971